OpenStudy (anonymous):

An exam consists of two parts, Section X and Section Y. There can be a maximum of 100 questions. There must be at least 5 more questions in Section Y than in Section X. Write a system of inequalities to model the number of questions in each of the two sections. Then solve the system by graphing.

OpenStudy (anonymous):

let there be x questions in section X, then there will be 100-x in section Y. And according to the question \[100-x \ge x+5\]

OpenStudy (anonymous):

then finally you get \[x \le 95/2 => x \le 47.5\]

OpenStudy (anonymous):